Find the lower limit on a 95% confidence interval for the height of men in the U. S. (Consider each man a sample of size 1.)
Mean = 70.21 inches
Std. Dev. = 1.02 inches

Answer by stanbon(57361)

Find the lower limit on a 95% confidence interval for the height of men in the U. S. (Consider each man a sample of size 1.)
Sample Mean = 70.21 inches
Sample Std. Dev. = 1.02 inches
------
ME = 1.96*1.02/sqrt(1) = 1.9992
------------------------
95% CI: 70.21-1.9992 < u < 70.21+1.9992
----
Lower Limit: 70.21-1.9992 = 68.21
====================================
Cheers,
Stan H.
===========